pandasurvive:

These are the values in Rita’s data set. (1, 19), (2, 33), (4,45), (5, 57), (6,71.5) Rita determines the equation of a linear regression line to be yˆ=9.7x+10.3 . Use the point tool to graph the residual plot for the data set. Round residuals to the nearest unit as needed.

pandasurvive:

1 attachment
Vocaloid:

hints: residual = observed - predicted use the equation provided yˆ=9.7x+10.3 to calculate the predicted y^ value for each x value then calculate the residuals then plot them on the graph

pandasurvive:

Yeah sorry @Vocaloid have to keep getting on and off the site.

pandasurvive:

19^=9.7(1)+10.3 is that how I would do it?

Vocaloid:

19 is your observed value, not predicted, so you would instead: y^ = 9.7(1) + 10.3 and solve for y^

pandasurvive:

ohhh, and do that for all my numbers?

Vocaloid:

yup

pandasurvive:

so then i would have to do 19-20=1 and so on?

Vocaloid:

*-1, yes. that would be your residual so you would plot (1,-1)

Vocaloid:

I've done the calculations already, so when you are done, you can check them against mine

pandasurvive:

(2,3.3) (4,4.1) (5,1.8) (6,3) is what i got

pandasurvive:

that was before i rounded though

Vocaloid:

check your signs, remember it is observed - predicted

Vocaloid:

some of your residuals should be negative

pandasurvive:

(2,3.3) (4,-4.1) (5-,1.8) (6,3) then they round to (2,3) (4,-4) (5,-2) (6,3)

Vocaloid:

perfect! now you can plot your residuals

pandasurvive:

Thanks :D

pandasurvive:

Sorry I took so long, I am not the best at understanding things, lol.

Vocaloid:

nah, you're good, don't apologize

pandasurvive:

Ok well thanks :D